OpenStudy (hanna12):

Can someone help me match the set of vertices with the type of quadrilateral they form? A(2, 0), B(3, 2), C(6, 3), D(5, 1) a parallelogram with nonperpendicular adjacent sides A(3, 3), B(3, 6), C(7, 6), D(7, 3) a rectangle with noncongruent adjacent sides A(-5, -3), B(-4, -1), C(-1, -1), D(0, -3) a square A(2, -2), B(3, 0), C(4, -2), D(3, -4) a rhombus with nonperpendicular adjacent sides A(3, 3), B(2, 5), C(4, 6), D(5, 4)

OpenStudy (luffingsails):

Try taking one of the sets of points and graphing them... See what you get and see which of the explanations of figures matches the set of points.

OpenStudy (luffingsails):

As an example consider the first set of points: A(2, 0), B(3, 2), C(6, 3), D(5, 1). When graphed it looks like this:

OpenStudy (luffingsails):

In this case, thanks to @mathmate, it is a Rhombus.

OpenStudy (mathstudent55):

It is not a rhombus because not all sides are congruent.

OpenStudy (mathmate):

Right, it's a parallelogram. I was looking at it in small scale and thought "it looks like a rhombus".

OpenStudy (mathmate):

If you prefer to work with the coordinates, you can do so using three simple tests, 1. to test if it is a parallelgram: check if \(x_a+x_c=x_b+x_d\) 2. to test if adjacent sides are equal (for squares and rhombuses) check if \((x_b−x_a)^2+(y_b−y_a)^2=(x_c−x_b)^2+(y_c−y_b)^2\) Note: you don't even need to square the numbers if the numbers are like \((−3)^2+4^2\) and \(4^2+3^2\) because we know that after squaring, the sums will be identical. 3. to test if adjacent sides are perpendicular (for squares, rectangles) check if \(\frac{yb−ya}{xb−xa}\times \frac{yc−yb}{xc−xb}=−1\) Once you have the information, you can determine which type of quadrilateral it is, among parallelogram, rhombus, square and rectangles. For examples and proofs, see: http://mathm8.altervista.org/geometry/index.html